Determine the boiling point of the following solutions: Kb(water) = 0.52 oC kg/mol ,0.90 m Sr(ClO3)2 Assume 100 % dissociation of Sr(ClO3)2 in solution
Answer #1

Sr(ClO3)2 --> Sr2+ + 2ClO3-

Given 100% dissociation so i = 3

Now, dT = i*kb*m

dT = 3*0.52*0.90 = 1.404 °C

so boiling point of water = 100 + 1.404 = 101.404°C

If you have any query please comment...

If satisfied please rate it thanks...